OpenStudy (sasogeek):

if 1+1=2 and 1-1=0 and 1+2+3+4+5=15 what is 5-4-3-2-1=?

OpenStudy (amistre64):

5 - 10

OpenStudy (chaise):

Don't you own a calculator? -5

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=5-4-3-2-1

OpenStudy (amistre64):

i own a mancala board

myininaya (myininaya):

if 1+2+3+4+5=15 then -(1+2+3+4+5)=-15 so -1-2-3-4-5=-15 add 10 on both sides -1-2-3-4-5+10=-15+10 -1-2-3-4+5=-5

myininaya (myininaya):

and of course we can use everything you have before what is so we have 1+1=2 5(1+1)=5(2) 5(1+1)=10

myininaya (myininaya):

so i added 5(1+1) on both sides

myininaya (myininaya):

but remember 1+1=2

myininaya (myininaya):

and 5(2)=10

myininaya (myininaya):

so we can say i added 5(1+1) on both sides or we can say i added 10 on both sides

myininaya (myininaya):

whatever floats your boat

OpenStudy (anonymous):

mancala lol

myininaya (myininaya):

lol

OpenStudy (anonymous):

i graduated to an abacus

OpenStudy (anonymous):

it is still beta, but works ok only a few bugs

myininaya (myininaya):

why does wolfram say minus 5 instead of negative 5

OpenStudy (sasogeek):

idk

OpenStudy (anonymous):

i too say "minus 5"

myininaya (myininaya):

negative 5 i believe is more correct

OpenStudy (anonymous):

use "negative" as an adjective. as in "that is a negative number"

myininaya (myininaya):

minus is an operation

OpenStudy (chaise):

Minus five and negative 5 is the same. When you say minus 5, it's assumed minus 5 from 0.

myininaya (myininaya):

i know 0 minus 5 and negative 5 are the same

OpenStudy (anonymous):

i love these kind of maths -5

OpenStudy (anonymous):

the sign comes with the number - 5 read "minus 5"

myininaya (myininaya):

0 minus 5 is right not minus 5 lol its my turn to pick on you satellite